If \\(\\bar{PQ}=4I + 3j + 2k\\), then find \\(|\\bar {PQ}|\\)
Answer by jim_thompson5910(35256)


PQ = 4i + 3j + 2k is in the form ai + bj + ck

The length of vector ai + bj + ck is sqrt(a^2 + b^2 + c^2), so

|PQ| = sqrt(4^2 + 3^2 + 2^2) = sqrt(16+9+4) = sqrt(29) = 5.38516

Exact Answer: sqrt(29)
Approximate Answer: 5.38516